OpenStudy (anonymous):

Write the expression as either the sine, cosine, or tangent of a single angle. (I will draw the problem as best as I can).

OpenStudy (anonymous):

\[\cos((\pi/3))\cos((\pi/5))+ \sin((\pi/3))\sin((\pi/5))\]